Using exponential smoothing of the data, forecast the sales for the month of July. actual sale: March 1,210 April 1,280, May 1,260 June 1,275

Forecast April 1,240

Alpha 0.20
Answer #1
Period Actual [A(t)]
March 1210
April 1280
May 1260
June 1275
F(t) = F(t-1) + α [A(t-1) - F(t-1)]
Period Exponential Forecast [F(t)]
F(t-1) + α [A(t-1) - F(t-1)]
April Given = 1240
May 1240 + 0.20 [1280 - 1240] = 1248
June 1248 + 0.20 [1260 - 1248] = 1250
July 1250 + 0.20 [1275 - 1250] = 1255
Therefore, As per the Exponential Forecast, Sales for July is 1255